OpenStudy (annmeez):

the area of triangle LMN is 18ft squared, and the area of triangle FGH is 24ft squared. If the two triangles are similar what is the ratio of LM to FG

OpenStudy (welshfella):

the ratio of corresponding sides = square root of ratio of theire areas

OpenStudy (welshfella):

so what is the square root of 18 squared, and of 24 squared?

OpenStudy (annmeez):

18 is 3 square root 2 and 24 is 2 square root 6 ?

OpenStudy (welshfella):

No i didn't ask for the square root of 18 but the square root of 18^2

OpenStudy (welshfella):

what number multiplied by itself gives 18^2?

OpenStudy (welshfella):

18^2 = 18 * 18 right?

OpenStudy (welshfella):

maybe its too obvious ...

OpenStudy (welshfella):

it' simpler than you realise

OpenStudy (welshfella):

what is square root of 4 Its 2 so 2 *2 = 4

OpenStudy (welshfella):

* as 2*2 = 4